OpenStudy (anonymous):

what number gives a result of -5 when 9 is subtracted from the quotient of the number and 6?

OpenStudy (mertsj):

The quotient of the number and 6 is x/6

OpenStudy (mertsj):

If you subtract 9 from that quotient, you would get -5 according to the problem. Write and solve the equation.

OpenStudy (anonymous):

the equation would be y-5=9-y/6 @Mertsj that's what I got

OpenStudy (mertsj):

\[\frac{x}{6}-9=-5\]